# DyGMamba: Efficiently Modeling Long-Term Temporal Dependency on Continuous-Time Dynamic Graphs with State Space Models

### v1

@@ 1 Introduction
Dynamic graphs store node interactions in the form of links labeled with timestamps (Kazemi et al., 2020). In recent years, learning dynamic graphs has gained increasing interest since it can be used to facilitate various real-world applications, e.g., social network analysis (Huo et al., 2018) and recommender system development (Wang et al., 2023). Dynamic graphs can be classified into two types, i.e., discrete-time dynamic graph (DTDG) and continuous-time dynamic graph (CTDG). A DTDG is represented as a sequence of graph snapshots that are observed at regular time intervals, where all the edges in a snapshot are taken as existing simultaneously, while a CTDG consists of a stream of events where each of them is observed individually with its own timestamp. Previous work (Kazemi et al., 2020) has indicated that CTDGs have an advantage over DTDGs in preserving temporal details, and therefore, more attention is paid to developing novel CTDG modeling approaches for dynamic graph representation learning.
@@
Recent effort in CTDG modeling has resulted in a wide range of models (Trivedi et al., 2019; Xu et al., 2020; Rossi et al., 2020; Cong et al., 2023). However, most of these methods are unable to model long-term temporal dependencies of nodes. To solve this problem, (Yu et al., 2023) proposes a CTDG model DyGFormer that can handle long-term node interaction histories based on Transformer (Vaswani et al., 2017). Despite its ability in modeling longer histories, employing a Transformer naturally introduces excessive usage of computational resources due to its quadratic complexity. Another recent work CTAN (Gravina et al., 2024) tries to capture long-term temporal dependencies by propagating graph information in a non-dissipative way over time with a graph convolution-based model. Despite the model’s high efficiency, (Gravina et al., 2024) shows that it cannot capture very long histories and is surpassed by (Yu et al., 2023) on the CTDGs where learning from very far away temporal information is critical. Based on these observations, we summarize the first challenge in CTDG modeling: How to develop a model that is scalable in modeling very long-term historical interactions? Another point worth noting is that as longer histories introduce more temporal information, more powerful models are needed to identify and select the most critical parts. This reveals another challenge: How to effectively select critical temporal information with long node interaction histories?
@@
To address the first challenge, we propose to leverage a currently popular state space model (SSM), i.e., Mamba SSM (Gu & Dao, 2023) to encode the long sequence of historical node interactions. Since Mamba is proven effective and efficient in long sequence modeling (Gu & Dao, 2023), it maintains low computational complexity and is scalable in modeling long-term temporal dependencies. For the second challenge, we address it by learning temporal patterns of node interactions and dynamically select the critical temporal information based on them. The motivation can be explained by the following example. Consider a CTDG with nodes as people or songs and edges representing a person playing a song at a specific time. If a person $u$ frequently plays a hit song $v$ initially but decreases the frequency as $v$’s popularity declines, the time intervals between plays increase. This increasing interval forms a temporal pattern for the edge $(u,v,t)$ between $u$ and $v$ at a future timestamp $t$. Ignoring this pattern can lead models to incorrectly predict that $u$ will still play $v$ at future timestamps due to their high appearances in each other’s historical interactions, resulting in poor performance. However, if a CTDG model is aware of this pattern, it can be a hint for the model to pay less attention to the interactions between $u$ and $v$, and select other part of temporal information which are more contributive in making correct predictions (e.g., $u$ starts to listen to another song $v^{\prime}$ a lot before $t$). Since each of such pattern corresponds to a specific edge, e.g., $(u,v,t)$, we name these patterns as edge-specific temporal patterns.
@@
To this end, we propose a new CTDG model named DyGMamba. DyGMamba first leverages a node-level Mamba SSM to encode historical node interactions. Another time-level Mamba SSM is then employed to exploit the edge-specific temporal patterns, where its output is used to dynamically select the critical information from the interaction history. To summarize: (1) We present DyGMamba, the first model using SSMs for CTDG representation learning; (2) DyGMamba demonstrates high efficiency in modeling long-term temporal dependencies in CTDGs; (3) Experimental results show that DyGMamba achieves new state-of-the-art on dynamic link prediction over most commonly-used CTDG datasets. It shows even stronger superiority as the dataset requires longer temporal dependency.
@@ 2.1 Related Work
Dynamic graph representation learning methods can be categorized into two groups, i.e., DTDG and CTDG methods. DTDG methods (Pareja et al., 2020; Goyal et al., 2020; Sankar et al., 2020; You et al., 2022; Li et al., 2024) can only model DTDGs where each of them is represented as a sequence of graph snapshots. Modeling a dynamic graph as graph snapshots require time discretization and will inevitably cause information loss (Kazemi et al., 2020). To overcome this problem, recent works focus more on developing CTDG methods that treat a dynamic graph as a stream of events, where each event has its own unique timestamp. Some works (Trivedi et al., 2019; Chang et al., 2020) model CTDGs by using temporal point process. Another line of works (Xu et al., 2020; Ma et al., 2020; Wang et al., 2021b; Gravina et al., 2024; Petrović et al., 2024) designs advanced temporal graph neural networks for CTDGs. Besides, some other methods are developed based on memory networks (Kumar et al., 2019; Rossi et al., 2020; Liu et al., 2022), temporal random walk (Wang et al., 2021c; Jin et al., 2022) and temporal sequence modeling (Wang et al., 2021a; Cong et al., 2023; Yu et al., 2023; Besta et al., 2023; Tian et al., 2024). Since some real-world CTDGs heavily rely on long-term temporal information for effective learning, a number of works start to study how to build CTDG models that can do long range propagation of information over time (Yu et al., 2023; Gravina et al., 2024). We compare them later in experiments with DyGMamba to show our model outperforms them in capturing long-term temporal information.
@@
Transformer (Vaswani et al., 2017) is a de facto backbone architecture in modern deep learning. However, its self-attention mechanism results in large space and time complexity, making it unsuitable for long sequence modeling (Duman Keles et al., 2023). To address this, many works focus on building structured state space models that scale linearly or near-linearly with input sequence length (Gu et al., 2021; 2022b; 2022a; Smith et al., 2023; Peng et al., 2023; Ma et al., 2023; Gu & Dao, 2023). Most structured SSMs exhibit linear time invariance (LTI), meaning their parameters are not input-dependent and fixed for all time-steps. (Gu & Dao, 2023) demonstrates that LTI prevents SSMs from effectively selecting relevant information from the input context, which is problematic for tasks requiring context-aware reasoning. To solve this issue, (Gu & Dao, 2023) proposes S6, also known as Mamba, which uses a selection mechanism to dynamically choose important information from input sequence elements. Selection mechanism involves learning functions that map input data to the SSM’s parameters, making Mamba both efficient and effective in modeling language, DNA sequences, and audio. Recently, there have been several works employing Mamba SSM for representation learning on static graphs (Wang et al., 2024; Behrouz & Hashemi, 2024), however, they are not designed for dynamic graphs. Another work (Li et al., 2024) conducts a preliminary study on DTDG modeling with SSMs but it cannot be generalized to CTDGs.
@@ 2.2 Preliminaries
A structured SSM (Gu et al., 2022b) is inspired by a continuous system described below:
@@
$q(\tau)\in\mathbb{R}$ and $r(\tau)\in\mathbb{R}$ are the 1-dimensional input and output over time $\tau$ [^1] We use $\tau$ rather than $t$ to indicate time in a a continuous system to distinguish from the time in CTDGs., respectively. $\mathbf{A}\in\mathbb{R}^{d_{1}\times d_{1}},\mathbf{B}\in\mathbb{R}^{1\times d_{1}},\mathbf{C}\in\mathbb{R}^{d_{1}\times 1}$ are three parameters deciding the system. Based on it, a structured SSM, i.e., S4 (Gu et al., 2022b), includes a time-scale parameter $\Delta\in\mathbb{R}$ and discretizes all the parameters to adapt the model to processing sequences recurrently.
@@ 3.1 Learning One-Hop Temporal Neighbors
Given one-hop temporal neighbors $\textit{Nei}_{u}^{t}$ of the source node $u$, we sort them in the chronological order and append $(u,t)$ at the end to form a sequence of $\textit{Nei}_{u}^{t}+1$ temporal nodes. We take their node features from the dataset and stack them into a feature matrix $\tilde{\mathbf{X}}_{u}^{t}\in\mathbb{R}^{(|\textit{Nei}_{u}^{t}|+1)\times d_{N}}$. Similarly, we build a link feature matrix $\tilde{\mathbf{E}}_{u}^{t}\in\mathbb{R}^{(|\textit{Nei}_{u}^{t}|+1)\times d_{E}}$. To incorporate temporal information, we encode the time difference between $u$ and each one-hop temporal neighbor $(u^{\prime},t^{\prime})$ using the time encoding function introduced in TGAT (Xu et al., 2020): $\sqrt{1/d_{T}}[cos(\omega_{1}(t-t^{\prime})+\phi_{1}),\dots,cos(\omega_{d}(t-t^{\prime})+\phi_{d_{T}})]$, where $d_{T}$ is the dimension of time representation, $\omega_{1}\dots\omega_{d_{T}}$ and $\phi_{1}\dots\phi_{d_{T}}$ are trainable parameters. The time feature of $u$’s temporal neighbors are denoted as $\tilde{\mathbf{T}}_{u}^{t}\in\mathbb{R}^{(|\textit{Nei}_{u}^{t}|+1)\times d_{T}}$. We follow the same way to get $\tilde{\mathbf{X}}_{v}^{t}\in\mathbb{R}^{(|\textit{Nei}_{v}^{t}|+1)\times d_{N}}$, $\tilde{\mathbf{E}}_{v}^{t}\in\mathbb{R}^{(|\textit{Nei}_{v}^{t}|+1)\times d_{E}}$ and $\mathbf{T}_{v}^{t}\in\mathbb{R}^{(|\textit{Nei}_{v}^{t}|+1)\times d_{T}}$ for $v$’s temporal neighbors. Following (Tian et al., 2024), we also consider the historical node interaction frequencies in the interaction sequences $\mathcal{S}_{u}^{t}$ and $\mathcal{S}_{v}^{t}$ of source $u$ and destination $v$. For example, assume the interacted nodes of $u$ and $v$ (arranged in chronological order) are $\{a,v,a\}$ and $\{b,b,u,a\}$, the appearing frequencies of $a$, $b$ in $u$/$v$’s historical interactions are 2/1, 0/2, respectively. And the frequency of the interaction involving $u$ and $v$ is 1. Thus, the node interaction frequency features of $u$ and $v$ are written as $\tilde{F}_{u}^{t}=[[2,1],[1,1],[2,1],[0,1]]^{\top}$ and $\tilde{F}_{v}^{t}=[[0,2],[0,2],[1,1],[2,1],[0,1]]^{\top}$, respectively. Note that the last elements ($[0,1]$ and $[0,1]$) in $\tilde{F}_{u}^{t}$ and $\tilde{F}_{v}^{t}$ correspond to the appended $(u,t)$ and $(v,t)$ not existing in the observed histories. We initialize them with [0,number of historical interactions $[0,\textit{number of historical interactions between}\ u,v]$. An encoding multilayer perceptron (MLP) $f(\cdot):\mathbb{R}\to\mathbb{R}^{d_{F}}$ is employed to encode these features into representations: $\tilde{\mathbf{F}}_{u}^{t}=f(\tilde{F}_{u}^{t}[:,0])+f(\tilde{F}_{u}^{t}[:,1])\in\mathbb{R}^{(|\textit{Nei}_{u}^{t}|+1)\times d_{F}}$, $\tilde{\mathbf{F}}_{v}^{t}=f(\tilde{F}_{v}^{t}[:,0])+f(\tilde{F}_{v}^{t}[:,1])\in\mathbb{R}^{(|\textit{Nei}_{v}^{t}|+1)\times d_{F}}$.
@@
We employ the patching technique proposed by (Yu et al., 2023) to save computational resources when dealing with a large number of temporal neighbors. We treat $p$ temporally adjacent neighbors as a patch and flatten their features. For example, with patching, $\tilde{\mathbf{X}}_{u}^{t}\in\mathbb{R}^{(|\textit{Nei}_{u}^{t}|+1)\times d_{N}}$ results in a new patched feature matrix $\mathbf{X}_{u}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{u}^{t}|+1)/p\rceil\times(p\cdot d_{N})}$ (we pad $\tilde{\mathbf{X}}_{u}^{t}$ with zero-valued features when $|\textit{Nei}_{u}^{t}|+1$ cannot be divided by $p$). Similarly, we get $\mathbf{X}_{\theta}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times(p\cdot d_{N})}$, $\mathbf{E}_{\theta}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times(p\cdot d_{E})}$, $\mathbf{T}_{\theta}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times(p\cdot d_{T})}$ and $\mathbf{F}_{\theta}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times(p\cdot d_{F})}$ ($\theta$ is either $u$ or $v$). Each row of a feature matrix (whether before or after patching) corresponds to an element of the input sequence sent into an SSM later. Recall that SSMs process sequences in a recurrent way. Patching decreases the length of the sequence by roughly $p$ times, making great contribution in saving computational resources.
@@
$\mathbf{W}_{\mathbf{B}_{1}},\mathbf{W}_{\mathbf{C}_{1}}\in\mathbb{R}^{4d\times d_{\text{SSM}}}$ and $\mathbf{W}_{\Delta_{1}}\in\mathbb{R}^{4d\times 1}$. $\bar{\mathbf{A}}_{1},\bar{\mathbf{B}}_{1}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times 4d\times d_{\text{SSM}}}$ are discretized parameters. $\text{Param}_{\Delta_{1}}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times 4d}$ is a parameter defined in (Gu & Dao, 2023). $\mathbf{I}$ is an identity matrix, and Eq. 4d denotes the SSM forward process similar to Eq. 2. Then we use an MLP $f_{\text{node}}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$ on SSM’s output
@@
After $l_{N}$ layers, we have $\mathbf{H}_{u}^{t}$ and $\mathbf{H}_{v}^{t}$ that contain the encoded information of all one-hop temporal neighbors for the entities $u$ and $v$ as well as the information of themselves. Since we sort temporal neighbors chronologically, our node-level SSM block can directly learn the temporal dynamics for graph forecasting. Another point worth noting is that Mamba SSM is originally good at capturing long-term dependencies. Our model can benefit from this when temporally faraway neigbors are critical for decision making.

### v2

@@ 1 Introduction
Dynamic graphs store node interactions in the form of links labeled with timestamps (Kazemi et al., 2020). In recent years, learning dynamic graphs has gained increasing interest since it can be used to facilitate various real-world applications. Dynamic graphs can be classified into two types, i.e., discrete-time dynamic graph (DTDG) and continuous-time dynamic graph (CTDG). A DTDG is represented as a sequence of graph snapshots that are observed at regular time intervals, where all the edges in a snapshot are taken as existing simultaneously, while a CTDG consists of a stream of events where each of them is observed individually with its own timestamp. Previous work (Kazemi et al., 2020) has indicated that CTDGs have an advantage over DTDGs in preserving temporal details, and therefore, more attention is paid to developing novel CTDG modeling approaches for dynamic graph representation learning.
@@
Recent effort in CTDG modeling has resulted in a wide range of models. However, most of them are unable to model long-term temporal dependencies of nodes. To solve this problem, Yu et al. (2023) propose a CTDG model DyGFormer that can handle long-term node interaction histories based on Transformer (Vaswani et al., 2017). Despite its ability in modeling longer histories, employing a Transformer naturally introduces excessive usage of computational resources due to its quadratic complexity. Another recent work CTAN (Gravina et al., 2024) tries to capture long-term temporal dependencies by propagating graph information in a non-dissipative way over time with a graph convolution-based model. Despite the model’s high efficiency, Gravina et al. (2024) show that CTAN cannot capture very long histories and is surpassed by DyGFormer on the CTDGs where learning from very far away temporal information is critical. Based on these observations, we summarize the first challenge in CTDG modeling: How to develop a model that is scalable in modeling very long-term historical interactions? Another point worth noting is that as longer histories introduce more temporal information, more powerful models are needed to identify and select the most critical parts. This reveals another challenge: How to effectively select critical temporal information with long node interaction histories?
@@
To address the first challenge, we propose to leverage a popular state space model (SSM), i.e., Mamba SSM (Gu & Dao, 2023) to encode the long sequence of historical node interactions. Since Mamba is proven effective and efficient in long sequence modeling (Gu & Dao, 2023), it maintains low computational complexity and is scalable in modeling long-term temporal dependencies. For the second challenge, we address it by learning temporal patterns of node interactions and dynamically selecting the critical temporal information based on them. The motivation can be explained by the following example. Consider a CTDG with nodes as people or songs and edges representing a person playing a song at a specific time. If a person $u$ frequently plays a hit song $v$ initially but decreases the frequency later on, the time intervals between plays increase. Ignoring this pattern can lead models to incorrectly predict that $u$ will still play $v$ at future timestamps due to their high appearances in each other’s historical interactions. If a CTDG model recognizes this pattern, it can prioritize other temporal information, such as $u$ increasingly listening to a new song $v^{\prime}$ before $t$, instead of focusing on $u$, $v$ interactions. Since each pattern corresponds to a specific edge, e.g., $(u,v,t)$, we name these patterns as edge-specific temporal patterns.
@@
To this end, we propose a new CTDG model named DyGMamba. DyGMamba first leverages a node-level Mamba SSM to encode historical node interactions. Another time-level Mamba SSM is then employed to exploit the edge-specific temporal patterns, where its output is used to dynamically select the critical information from the interaction history. To summarize: (1) We present DyGMamba, the first model using SSMs for CTDG representation learning; (2) DyGMamba demonstrates high efficiency and strong efectiveness in modeling long-term temporal dependencies in CTDGs; (3) Experimental results show that DyGMamba achieves new state-of-the-art on dynamic link prediction over most commonly-used CTDG datasets.
@@ 2.1 Related Work
Dynamic graph representation learning methods can be categorized into two groups, i.e., DTDG and CTDG methods. DTDG methods (Pareja et al., 2020; Goyal et al., 2020; Sankar et al., 2020; You et al., 2022; Li et al., 2024) can only model DTDGs where each of them is represented as a sequence of graph snapshots. Modeling a dynamic graph as graph snapshots requires time discretization and will inevitably cause information loss (Kazemi et al., 2020). To overcome this problem, recent works focus more on developing CTDG methods that treat a dynamic graph as a stream of events, where each event has its own unique timestamp. Some works (Trivedi et al., 2019; Chang et al., 2020) model CTDGs by using temporal point process. Another line of works (Xu et al., 2020; Ma et al., 2020; Wang et al., 2021b; Gravina et al., 2024) designs advanced temporal graph neural networks for CTDGs. Besides, some other methods are developed based on memory networks (Kumar et al., 2019; Rossi et al., 2020; Liu et al., 2022), temporal random walk (Wang et al., 2021c; Jin et al., 2022) and temporal sequence modeling (Cong et al., 2023; Yu et al., 2023; Besta et al., 2023; Tian et al., 2024). Since some real-world CTDGs heavily rely on long-term temporal information for effective learning, a number of works start to study how to build CTDG models that can do long range propagation of information over time (Yu et al., 2023; Gravina et al., 2024). We compare them later in experiments with DyGMamba to show our model outperforms them in capturing long-term temporal information.
@@
Transformer (Vaswani et al., 2017) is a de facto backbone architecture in modern deep learning. However, its self-attention mechanism results in large space and time complexity, making it unsuitable for extremely long sequence modeling (Duman Keles et al., 2023). To address this, many works focus on building structured state space models that scale linearly or near-linearly with input sequence length (Gu et al., 2021; 2022b; 2022a; Smith et al., 2023; Peng et al., 2023; Ma et al., 2023; Gu & Dao, 2023). Most structured SSMs exhibit linear time invariance (LTI), meaning their parameters are not input-dependent and fixed for all time-steps. Gu & Dao (2023) demonstrate that LTI prevents SSMs from effectively selecting relevant information from the input context, which is problematic for tasks requiring context-aware reasoning. To solve this issue, Gu & Dao (2023) proposes S6, also known as Mamba, which uses a selection mechanism to dynamically choose important information from input sequence elements. Selection mechanism involves learning functions that map input data to SSM’s parameters, making Mamba both efficient and effective in modeling language, DNA sequences, and audio. In our work, we show how SSM can be used to achieve effective and efficient representation learning on CTDGs.
@@ 2.2 Preliminaries
A structured SSM, i.e., S4 (Gu et al., 2022b), is inspired by a continuous system described below:
@@
$q(\tau)\in\mathbb{R}$ and $r(\tau)\in\mathbb{R}$ are the 1-dimensional input and output over time $\tau$ [^1] We use $\tau$ rather than $t$ to indicate time in a continuous system to distinguish from the time in CTDGs., respectively. $\mathbf{A}\in\mathbb{R}^{d_{1}\times d_{1}},\mathbf{B}\in\mathbb{R}^{1\times d_{1}},\mathbf{C}\in\mathbb{R}^{d_{1}\times 1}$ are three parameters deciding the system. Based on it, an S4 includes a time-scale parameter $\Delta\in\mathbb{R}$ and discretizes all the parameters to adapt the model to processing sequences recurrently.
@@ 3.1 Learning One-Hop Temporal Neighbors
Given one-hop temporal neighbors $\textit{Nei}_{u}^{t}$ of the source node $u$, we sort them in the chronological order and append $(u,t)$ at the end to form a sequence of $\textit{Nei}_{u}^{t}+1$ temporal nodes. We take their node features from the dataset and stack them into a feature matrix $\tilde{\mathbf{X}}_{u}^{t}\in\mathbb{R}^{(|\textit{Nei}_{u}^{t}|+1)\times d_{N}}$. Similarly, we build a link feature matrix $\tilde{\mathbf{E}}_{u}^{t}\in\mathbb{R}^{(|\textit{Nei}_{u}^{t}|+1)\times d_{E}}$. To incorporate temporal information, we encode the time difference between $u$ and each one-hop temporal neighbor $(u^{\prime},t^{\prime})$ using the time encoding function introduced in TGAT (Xu et al., 2020): $\sqrt{1/d_{T}}[cos(\omega_{1}(t-t^{\prime})+\phi_{1}),\dots,cos(\omega_{d}(t-t^{\prime})+\phi_{d_{T}})]$. $d_{T}$ is the dimension of time representation. $\omega_{1}\dots\omega_{d_{T}}$ and $\phi_{1}\dots\phi_{d_{T}}$ are trainable parameters. The time feature of $u$’s temporal neighbors are denoted as $\tilde{\mathbf{T}}_{u}^{t}\in\mathbb{R}^{(|\textit{Nei}_{u}^{t}|+1)\times d_{T}}$. We follow the same way to get $\tilde{\mathbf{X}}_{v}^{t}\in\mathbb{R}^{(|\textit{Nei}_{v}^{t}|+1)\times d_{N}}$, $\tilde{\mathbf{E}}_{v}^{t}\in\mathbb{R}^{(|\textit{Nei}_{v}^{t}|+1)\times d_{E}}$ and $\tilde{\mathbf{T}}_{v}^{t}\in\mathbb{R}^{(|\textit{Nei}_{v}^{t}|+1)\times d_{T}}$ for $v$’s temporal neighbors. Following Tian et al. (2024), we also consider the historical node interaction frequencies in the interaction sequences $\mathcal{S}_{u}^{t}$ and $\mathcal{S}_{v}^{t}$ of source $u$ and destination $v$. For example, assume the interacted nodes of $u$ and $v$ (arranged in chronological order) are $\{a,v,a\}$ and $\{b,b,u,a\}$, the appearing frequencies of $a$, $b$ in $u$/$v$’s historical interactions are 2/1, 0/2, respectively. And the frequency of the interaction involving $u$ and $v$ is 1. Thus, the node interaction frequency features of $u$ and $v$ are written as $\tilde{F}_{u}^{t}=[[2,1],[1,1],[2,1],[0,1]]^{\top}$ and $\tilde{F}_{v}^{t}=[[0,2],[0,2],[1,1],[2,1],[0,1]]^{\top}$, respectively. Note that the last elements ($[0,1]$ and $[0,1]$) in $\tilde{F}_{u}^{t}$ and $\tilde{F}_{v}^{t}$ correspond to the appended $(u,t)$ and $(v,t)$ not existing in the observed histories. We initialize them with [0,number of historical interactions $[0,\textit{number of historical interactions between}\ u,v]$. An encoding multilayer perceptron (MLP) $f(\cdot):\mathbb{R}\to\mathbb{R}^{d_{F}}$ is employed to encode these features into representations: $\tilde{\mathbf{F}}_{u}^{t}=f(\tilde{F}_{u}^{t}[:,0])+f(\tilde{F}_{u}^{t}[:,1])\in\mathbb{R}^{(|\textit{Nei}_{u}^{t}|+1)\times d_{F}}$, $\tilde{\mathbf{F}}_{v}^{t}=f(\tilde{F}_{v}^{t}[:,0])+f(\tilde{F}_{v}^{t}[:,1])\in\mathbb{R}^{(|\textit{Nei}_{v}^{t}|+1)\times d_{F}}$.
@@
We employ the patching technique proposed by (Yu et al., 2023) to save computational resources when dealing with a large number of temporal neighbors. We treat $p$ temporally adjacent neighbors as a patch and flatten their features. For example, with patching, $\tilde{\mathbf{X}}_{u}^{t}\in\mathbb{R}^{(|\textit{Nei}_{u}^{t}|+1)\times d_{N}}$ results in a new patched feature matrix $\mathbf{X}_{u}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{u}^{t}|+1)/p\rceil\times(p\cdot d_{N})}$ (we pad $\tilde{\mathbf{X}}_{u}^{t}$ with zero-valued features when $|\textit{Nei}_{u}^{t}|+1$ cannot be divided by $p$). Similarly, we get $\mathbf{E}_{\theta}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times(p\cdot d_{E})}$, $\mathbf{T}_{\theta}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times(p\cdot d_{T})}$ and $\mathbf{F}_{\theta}^{t}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times(p\cdot d_{F})}$ ($\theta$ is either $u$ or $v$). Each row of a feature matrix corresponds to an element of the input sequence sent into an SSM later. Recall that SSMs process sequences in a recurrent way. Patching decreases the length of the sequence by roughly $p$ times, making great contribution in saving computational resources.
@@
$\mathbf{W}_{\mathbf{B}_{1}},\mathbf{W}_{\mathbf{C}_{1}}\in\mathbb{R}^{4d\times d_{\text{SSM}}}$ and $\mathbf{W}_{\Delta_{1}}\in\mathbb{R}^{4d\times 1}$. $\bar{\mathbf{A}}_{1},\bar{\mathbf{B}}_{1}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times 4d\times d_{\text{SSM}}}$ are discretized parameters. $\text{Param}_{\Delta_{1}}\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil\times 4d}$ is a parameter defined by Gu & Dao (2023). $\mathbf{I}$ is an identity matrix, and Eq. 4d denotes the SSM forward process similar to Eq. 2. Then we use an MLP $f_{\text{node}}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$ on SSM’s output
@@
After $l_{N}$ layers, we have $\mathbf{H}_{u}^{t}$ and $\mathbf{H}_{v}^{t}$ that contain the encoded information of all one-hop temporal neighbors for the entities $u$ and $v$ as well as the information of themselves. Since we sort temporal neighbors chronologically, our node-level SSM block can directly learn the temporal dynamics for graph forecasting.